# thanks: Corresponding Author

### v1

@@ IV.2 $\mathcal{M}\rightarrow D^{\ast 0}\overline{D}^{\ast 0}$ and $D^{\ast+}D^{\ast-}$
$\widetilde{\Pi}_{\mu\nu}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{D^{\ast 0}}(y)$
×$\times J_{\nu}^{\overline{D}^{\ast 0}}(0)J^{\dagger}(x)\}|0\rangle.$ (65)
The correlation function $\widetilde{\Pi}_{\mu\nu}(p,p^{\prime})$ contains the interpolating currents $J_{\mu}^{D^{\ast 0}}(x)$ and $J_{\nu}^{\overline{D}^{\ast 0}}(x)$ of the vector mesons $D^{\ast 0}$ and $\overline{D}^{\ast 0}$, respectively:
@@
In terms of the parameters of the ground-level particles $\widetilde{\Pi}_{\mu\nu}(p,p^{\prime})$ is given by the expression
$\widetilde{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J_{\mu}^{D^{\ast 0}}|D^{\ast 0}(p^{\prime},\varepsilon_{1})\rangle}{p^{\prime 2}-m_{D^{\ast 0}}^{2}}\frac{\langle 0|J_{\nu}^{\overline{D}^{\ast 0}}|\overline{D}^{\ast 0}(q,\varepsilon_{2})\rangle}{q^{2}-m_{D^{\ast 0}}^{2}}$
×$\times\langle D^{\ast 0}(p^{\prime},\varepsilon_{1})\overline{D}^{\ast 0}(q,\varepsilon_{2})|\mathcal{M}(p)\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
@@
×$\left.\times\varepsilon_{1}^{\ast}\cdot\varepsilon_{2}^{\ast}-q\cdot\varepsilon_{1}^{\ast}p^{\prime}\cdot\varepsilon_{2}^{\ast}\right].$ (69)
Then $\Pi_{\mu\nu}(p,p^{\prime})$ expressed using parameters of the hadronic molecule $\mathcal{M}$, and mesons $D^{\ast 0}$, and $\overline{D}^{\ast 0}$ is given by the formula
$\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\widetilde{g}_{2}(q^{2})\Lambda f_{D^{\ast}}^{2}m_{D^{\ast 0}}^{2}}{\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{D^{\ast 0}}^{2}\right)(q^{2}-m_{D^{\ast 0}}^{2})}$
×$\times\left[\frac{1}{2}\left(m^{2}-m_{D^{\ast 0}}^{2}-q^{2}\right)g_{\mu\nu}-q_{\mu}p_{\nu}^{\prime}\right]+\cdots.$ (70)
The correlator $\Pi_{\mu\nu}(p,p^{\prime})$ computed by employing the quark propagators reads
$\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{b}b\rangle$
×$\times\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{ja}(y-x){}S_{c}^{ai}(x)\gamma_{\nu}S_{u}^{ij}(-y)\right].$ (71)
The correlator $\Pi_{\mu\nu}(p,p^{\prime})$ in both cases contains the two Lorentz structures $g_{\mu\nu}$ and $q_{\mu}p_{\nu}^{\prime}$. We use invariant amplitudes that correspond to structures proportional to $g_{\mu\nu}$ and derive a required sum rule.
@@
The decay constant $f_{D^{\ast}}$ of the mesons $D^{\ast 0}$ is fixed as $(252.2\pm 22.66)~{}\mathrm{MeV}$ The extrapolating function $\widetilde{\mathcal{G}}_{2}(Q^{2})$ has the parameters: $\widetilde{\mathcal{G}}_{2}^{0}=0.022~{}\mathrm{GeV}^{-1}$, $\widetilde{a}_{2}^{1}=6.70$, and $\widetilde{a}_{2}^{2}=-8.18$ (see, Fig. 4 ). Then, the strong coupling $\widetilde{g}_{2}$ amounts to
@@ IV.3 $\mathcal{M}\rightarrow D_{s}^{(\ast)+}D_{s}^{(\ast)-}$
The correlation function required to estimate the strong coupling $G_{1}$at the vertex $\mathcal{M}B^{+}B^{-}$ is given by the formula
$\widehat{\Pi}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J^{B^{-}}(y)$
×$\times J^{B^{+}}(0)J^{\dagger}(x)\}|0\rangle,$ (81)
@@
The contribution of the ground-state particles $\mathcal{M}$, $B^{+}$, and $B^{-}$ to $\widehat{\Pi}(p,p^{\prime})$ is
$\widehat{\Pi}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J^{B^{-}}|B^{-}(p^{\prime})\rangle}{p^{\prime 2}-m_{B}^{2}}\frac{\langle 0|J^{B^{+}}|B^{+}(q)\rangle}{q^{2}-m_{B}^{2}}$
×$\times\langle B^{-}(p^{\prime})B^{+}(q)|\mathcal{M}(p)\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots.$
To compute the phenomenological side of the SR $\widehat{\Pi}^{\mathrm{Phys}}(p,p^{\prime})$, we make use of the matrix elements
@@
Then for $\widehat{\Pi}^{\mathrm{Phys}}(p,p^{\prime})$, we get
$\widehat{\Pi}^{\mathrm{Phys}}(p,p^{\prime})=\frac{G_{1}(q^{2})\Lambda f_{B}^{2}m_{B}^{4}}{m_{b}^{2}\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{B}^{2}\right)}$
×$\times\frac{m^{2}+m_{B}^{2}-q^{2}}{2(q^{2}-m_{B}^{2})}+\cdots.$ (86)
The correlation function $\widehat{\Pi}(p,p^{\prime})$ in terms of the quark propagators reads
$\widehat{\Pi}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle$
×$\times\mathrm{Tr}\left[\gamma_{5}S_{b}^{ja}(y-x){}S_{b}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right].$ (87)
The heavy quark condensate $\langle\overline{c}c\rangle$ with some accuracy can be replaced by
@@
$\widehat{\Pi}^{\mathrm{OPE}}(p,p^{\prime})=-\frac{1}{36m_{c}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}$
×$\times\mathrm{Tr}\left[\gamma_{5}S_{b}^{ja}(y-x){}S_{b}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right].$ (89)
Numerical computations of the form factor $G_{1}(q^{2})$ are performed for -- $q^{2}=Q^{2}=2-10~{}\mathrm{GeV}^{2}$. The coupling $G_{1}$ is extracted at the $Q^{2}=-m_{B}^{2}$ by utilizing the fit function $\widehat{\mathcal{G}}_{1}(Q^{2})$ with parameters $\widehat{\mathcal{G}}_{1}^{0}=0.045~{}\mathrm{GeV}^{-1}$, $\widehat{a}_{2}^{1}=2.82$, and $\widehat{a}_{4}^{2}=-2.61$
@@ VI Concluding remarks
In this article, we have studied the scalar hadronic molecule $\mathcal{M}=B_{c}^{+}B_{c}^{-}$ in the context of QCD sum rule method. Our prediction for the mass of this molecule $m=(12725\pm 85)~{}\mathrm{MeV}$ provides important information about this state. It turns out that the molecular structure $\mathcal{M}$ is unstable against two-meson fall-apart processes.
@@
The fully heavy hadronic molecules $B_{c}^{(\ast)+}B_{c}^{(\ast)-}$ were explored in the framework of the coupled-channel unitary approach in Ref. Liu:2023gla . Results in this method depend on the parameter $\Lambda$ used to regularize some integrals. In the case of the scalar structure $B_{c}^{+}B_{c}^{-}$ at $\Lambda=600~{}\mathrm{MeV}$ the authors found the pole position $E_{\mathrm{p}}=(12503.3-i126.8)~{}\mathrm{MeV}$. In other words, the mass of the hadronic molecule $B_{c}^{+}B_{c}^{-}$ is $12503.3~{}\mathrm{MeV}$ which is smaller than one obtained in the present work. Nevertheless,  the molecule $B_{c}^{+}B_{c}^{-}$ in this picture is also unstable and can strongly decay into a pair of mesons $\eta_{b}\eta_{c}$.
@@
The mass of $\mathcal{M}$ has enabled us to fix its possible decay channels. We have computed partial widths of the two-meson dissociations $\mathcal{M}\rightarrow\eta_{b}\eta_{c},\ J/\psi\Upsilon$, and $B_{c}^{+}B_{c}^{-}$, which are dominant decay modes of the molecule $\mathcal{M}$. The annihilations of the $b\overline{b}$ and $c\overline{c}$ pairs in $\mathcal{M}$ give rise to other two-meson strong decays: We have considered numerous processes generated by this second mechanism. To calculate the widths of two-meson decays, we have employed the methods of the three-point sum rule approach. This is necessary to estimate strong couplings of particles at various $\mathcal{M}$-meson-meson vertices. The full width of $\mathcal{M}$ saturated by these decay modes amounts to $\Gamma\left[\mathcal{M}\right]=(155\pm 23)~{}\mathrm{MeV}$: The fall-apart processes constitute $66\%$ of this parameter.
@@
In our previous publications Agaev:2024wvp ; Agaev:2024mng ; Agaev:2025 , we performed comprehensive investigations of the four-quark mesons $0^{+}$, $1^{+}$ and $2^{+}$ with the quark content $bc\overline{b}\overline{c}$.  These states were treated as diquark-antidiquark structures. The scalar particle $T_{\mathrm{bc\overline{b}\overline{c}}}$ was modeled as a tetraquark composed of the scalar diquark and antidiquark components Agaev:2024wvp . The mass and full width of this tetraquark were found equal to $m=(12697\pm 90)~{}\mathrm{MeV}$ and $\Gamma\left[T_{\mathrm{bc\overline{b}\overline{c}}}\right]=(142.4\pm 16.9)~{}\mathrm{MeV}$, respectively.
@@
Comparing the masses of these two exotic structures, we see that the hadronic molecule $\mathcal{M}$ is heavier than the diquark-antidiquark state though differences between their masses and widths are small: In fact, within errors of computations, one may state that these two scalar compounds have approximately same masses. Their widths are also comparable with each other.

### v2

@@ IV.2 $\mathcal{M}\rightarrow D^{\ast 0}\overline{D}^{\ast 0}$ and $D^{\ast+}D^{\ast-}$
$\widetilde{\Pi}_{2\mu\nu}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{D^{\ast 0}}(y)$
$\times J_{\nu}^{\overline{D}^{\ast 0}}(0)J^{\dagger}(x)\}|0\rangle.$ (65)
The correlation function $\widetilde{\Pi}_{2\mu\nu}(p,p^{\prime})$ contains the interpolating currents $J_{\mu}^{D^{\ast 0}}(x)$ and $J_{\nu}^{\overline{D}^{\ast 0}}(x)$ of the vector mesons $D^{\ast 0}$ and $\overline{D}^{\ast 0}$, respectively:
@@
In terms of the parameters of the ground-level particles $\widetilde{\Pi}_{2\mu\nu}(p,p^{\prime})$ is given by the expression
$\widetilde{\Pi}_{2\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J_{\mu}^{D^{\ast 0}}|D^{\ast 0}(p^{\prime},\varepsilon_{1})\rangle}{p^{\prime 2}-m_{D^{\ast 0}}^{2}}\frac{\langle 0|J_{\nu}^{\overline{D}^{\ast 0}}|\overline{D}^{\ast 0}(q,\varepsilon_{2})\rangle}{q^{2}-m_{D^{\ast 0}}^{2}}$
$\times\langle D^{\ast 0}(p^{\prime},\varepsilon_{1})\overline{D}^{\ast 0}(q,\varepsilon_{2})|\mathcal{M}(p)\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
@@
$\left.\times\varepsilon_{1}^{\ast}\cdot\varepsilon_{2}^{\ast}-q\cdot\varepsilon_{1}^{\ast}p^{\prime}\cdot\varepsilon_{2}^{\ast}\right].$ (69)
Then $\widetilde{\Pi}_{2\mu\nu}(p,p^{\prime})$ expressed using parameters of the hadronic molecule $\mathcal{M}$, and mesons $D^{\ast 0}$, and $\overline{D}^{\ast 0}$ is given by the formula
$\widetilde{\Pi}_{2\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\widetilde{g}_{2}(q^{2})\Lambda f_{D^{\ast}}^{2}m_{D^{\ast 0}}^{2}}{\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{D^{\ast 0}}^{2}\right)(q^{2}-m_{D^{\ast 0}}^{2})}$
$\times\left[\frac{1}{2}\left(m^{2}-m_{D^{\ast 0}}^{2}-q^{2}\right)g_{\mu\nu}-q_{\mu}p_{\nu}^{\prime}\right]+\cdots.$ (70)
The correlator $\widetilde{\Pi}_{2\mu\nu}(p,p^{\prime})$ computed by employing the quark propagators reads
$\widetilde{\Pi}_{2\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{b}b\rangle$
$\times\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{ja}(y-x){}S_{c}^{ai}(x)\gamma_{\nu}S_{u}^{ij}(-y)\right].$ (71)
The correlator $\widetilde{\Pi}_{2\mu\nu}(p,p^{\prime})$ in both cases contains the two Lorentz structures $g_{\mu\nu}$ and $q_{\mu}p_{\nu}^{\prime}$. We use invariant amplitudes that correspond to structures proportional to $g_{\mu\nu}$ and derive a required sum rule.
@@
The decay constant $f_{D^{\ast}}$ of the mesons $D^{\ast 0}$ is fixed as $(252.2\pm 22.66)~\mathrm{MeV}$ . The extrapolating function $\widetilde{\mathcal{G}}_{2}(Q^{2})$ has the parameters: $\widetilde{\mathcal{G}}_{2}^{0}=0.022~\mathrm{GeV}^{-1}$, $\widetilde{a}_{2}^{1}=6.70$, and $\widetilde{a}_{2}^{2}=-8.18$ (see, Fig. 4 ). Then, the strong coupling $\widetilde{g}_{2}$ amounts to
@@ IV.3 $\mathcal{M}\rightarrow D_{s}^{(\ast)+}D_{s}^{(\ast)-}$
The correlation function required to estimate the strong coupling $G_{1}\$at the vertex $\mathcal{M}B^{+}B^{-}$ is given by the formula
$\widehat{\Pi}_{1}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J^{B^{-}}(y)$
$\times J^{B^{+}}(0)J^{\dagger}(x)\}|0\rangle,$ (81)
@@
The contribution of the ground-state particles $\mathcal{M}$, $B^{+}$, and $B^{-}$ to $\widehat{\Pi}_{1}(p,p^{\prime})$ is
$\widehat{\Pi}_{1}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J^{B^{-}}|B^{-}(p^{\prime})\rangle}{p^{\prime 2}-m_{B}^{2}}\frac{\langle 0|J^{B^{+}}|B^{+}(q)\rangle}{q^{2}-m_{B}^{2}}$
$\times\langle B^{-}(p^{\prime})B^{+}(q)|\mathcal{M}(p)\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots.$
To compute the phenomenological side of the SR $\widehat{\Pi}_{1}^{\mathrm{Phys}}(p,p^{\prime})$, we make use of the matrix elements
@@
Then for $\widehat{\Pi}_{1}^{\mathrm{Phys}}(p,p^{\prime})$, we get
$\widehat{\Pi}_{1}^{\mathrm{Phys}}(p,p^{\prime})=\frac{G_{1}(q^{2})\Lambda f_{B}^{2}m_{B}^{4}}{m_{b}^{2}\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{B}^{2}\right)}$
$\times\frac{m^{2}+m_{B}^{2}-q^{2}}{2(q^{2}-m_{B}^{2})}+\cdots.$ (86)
The correlation function $\widehat{\Pi}_{1}(p,p^{\prime})$ in terms of the quark propagators reads
$\widehat{\Pi}_{1}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle$
$\times\mathrm{Tr}\left[\gamma_{5}S_{b}^{ja}(y-x){}S_{b}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right].$ (87)
The heavy quark condensate $\langle\overline{c}c\rangle$ with some accuracy can be replaced by Shifman:1978bx
@@
$\widehat{\Pi}_{1}^{\mathrm{OPE}}(p,p^{\prime})=-\frac{1}{36m_{c}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}$
$\times\mathrm{Tr}\left[\gamma_{5}S_{b}^{ja}(y-x){}S_{b}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right].$ (89)
Numerical computations of the form factor $G_{1}(q^{2})$ are performed for - $q^{2}=Q^{2}=2-10~\mathrm{GeV}^{2}$. The coupling $G_{1}$ is extracted at the $Q^{2}=-m_{B}^{2}$ by utilizing the fit function $\widehat{\mathcal{G}}_{1}(Q^{2})$ with parameters $\widehat{\mathcal{G}}_{1}^{0}=0.045~\mathrm{GeV}^{-1}$, $\widehat{a}_{2}^{1}=2.82$, and $\widehat{a}_{4}^{2}=-2.61$
@@ VI Concluding remarks
In this article, we have studied the scalar hadronic molecule $\mathcal{M}=B_{c}^{+}B_{c}^{-}$ in the context of the QCD sum rule method. Our prediction for the mass of this molecule $m=(12725\pm 85)~\mathrm{MeV}$ provides important information about this state. It turns out that the molecular structure $\mathcal{M}$ is unstable against two-meson fall-apart processes.
@@
The fully heavy hadronic molecules $B_{c}^{(\ast)+}B_{c}^{(\ast)-}$ were explored in the framework of the coupled-channel unitary approach in Ref. Liu:2023gla . Results in this method depend on the parameter $\Lambda$ used to regularize some integrals. In the case of the scalar structure $B_{c}^{+}B_{c}^{-}$ at $\Lambda=600~\mathrm{MeV}$, the authors found the pole position $E_{\mathrm{p}}=(12503.3-i126.8)~\mathrm{MeV}$. In other words, the mass of the hadronic molecule $B_{c}^{+}B_{c}^{-}$ is $12503.3~\mathrm{MeV}$, which is smaller than the one obtained in the present work. Nevertheless,  the molecule $B_{c}^{+}B_{c}^{-}$ in this picture is also unstable and can decay strongly into the pair of mesons $\eta_{b}\eta_{c}$.
@@
The mass of $\mathcal{M}$ has enabled us to fix its possible decay channels. We have computed partial widths of the two-meson dissociations $\mathcal{M}\rightarrow\eta_{b}\eta_{c},\ J/\psi\Upsilon$, and $B_{c}^{+}B_{c}^{-}$, which are dominant decay modes of the molecule $\mathcal{M}$. The annihilations of the $b\overline{b}$ and $c\overline{c}$ pairs in $\mathcal{M}$ give rise to other two-meson strong decays: We have considered numerous processes generated by this second mechanism. To calculate the widths of two-meson decays, we have employed the method of the three-point sum rule. This is necessary to estimate strong couplings of particles at various $\mathcal{M}$-meson-meson vertices. The full width of $\mathcal{M}$ saturated by these decay modes amounts to $\Gamma\left[\mathcal{M}\right]=(155\pm 23)~\mathrm{MeV}$: The fall-apart processes constitute $66\%$ of this parameter.
@@
In our previous publications Agaev:2024wvp ; Agaev:2024mng ; Agaev:2025 , we performed comprehensive investigations of the four-quark mesons $0^{+}$, $1^{+}$ and $2^{+}$ with the quark content $bc\overline{b}\overline{c}$.  These states were treated as diquark-antidiquark structures. The scalar particle $T_{\mathrm{bc\overline{b}\overline{c}}}$ was modeled as a tetraquark composed of the scalar diquark and antidiquark Agaev:2024wvp . The mass and full width of this tetraquark were found equal to $m=(12697\pm 90)~\mathrm{MeV}$ and $\Gamma\left[T_{\mathrm{bc\overline{b}\overline{c}}}\right]=(142.4\pm 16.9)~\mathrm{MeV}$, respectively.
@@
Comparing the masses of these two exotic structures, we see that the hadronic molecule $\mathcal{M}$ is heavier than the diquark-antidiquark state, though differences between their masses and widths are small: In fact, within the errors of computations, one may state that these two scalar compounds have approximately the same masses. Their full widths are also comparable with each other, though there are identical decay channels with considerably different partial widths. For instance, the partial width of the decays into $B_{c}^{+}B_{c}^{-}$ mesons in the case of the diquark-antidiquark and molecule structures are equal to $33.9~\mathrm{MeV}$ and $17.7~\mathrm{MeV}$, respectively. Even lower and upper limits for these parameters, obtained by taking into account the errors of calculations, generate only negligible small overlapping regions. Therefore, such decay modes may be used to distinguish the molecular structure of the exotic state $bc\overline{b}\overline{c}$ from its diquark organization.